OpenStudy (anonymous):

2 and one-third plus 3 and one-half is equal to what number?

OpenStudy (anonymous):

2 1/3 + 3 1/2 = ? 7/3 + 7/2 = ? Does that make sense? They are equivalent

OpenStudy (anonymous):

You can then use cross multiplication to make the denominator the same. 14/6 + 21/6 = 35/6 Is that clear?